OpenStudy (anonymous):

64x^3+8=0 I know what the answers SHOULD be, I'm just so confused how to get there.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

are you asked to solve for x or expand?

OpenStudy (anonymous):

isn't it the sum of a square but in cubic form so you have 4x and 2 and then you get (4x+2) = 0 and (4x^2+8x+4)

OpenStudy (anonymous):

i'm having trouble when using the quadratic formula because i get that x = -1/2 (right) and x=-1 which is wrong.....

OpenStudy (anonymous):

you're just asked to solve for x it looks like.. that's pretty easy

OpenStudy (anonymous):

subtract 8 from both sides and divide both sides by 64

OpenStudy (anonymous):

x^3 = -1/8

OpenStudy (anonymous):

take the cube root of both sides x=-1/2

OpenStudy (anonymous):

i already got that part, i need the other two x's